# Open-Set Deepfake Detection: A Parameter-Efficient Adaptation Method with Forgery Style Mixture

### v2

@@ Abstract
Open-set face forgery detection poses significant security threats and presents substantial challenges for existing detection models. These detectors primarily have two limitations: they cannot generalize across unknown forgery domains and inefficiently adapt to new data. To address these issues, we introduce an approach that is both general and parameter-efficient for face forgery detection. Our method builds on the assumption that different forgery source domains exhibit distinct style statistics. Specifically, we design a forgery-style mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains. In addition, previous methods typically require fully fine-tuning pretrained networks, consuming substantial time and computational resources. Drawing on recent advancements in vision transformers (ViT) for face forgery detection, we develop a parameter-efficient ViT-based detection model that includes lightweight forgery feature extraction modules and enables the model to extract global and local forgery clues simultaneously. We only optimize the inserted lightweight modules during training, maintaining the original ViT structure with its pre-trained ImageNet weights. This training strategy effectively preserves the informative pre-trained knowledge while flexibly adapting the model to the task of Deepfake detection. Extensive experimental results demonstrate that the designed model achieves state-of-the-art generalizability with significantly reduced trainable parameters, representing an important step toward open-set Deepfake detection in the wild.
@@ I Introduction
Face forgery detection has long posed a challenging yet crucial problem. Early methodologies proposed extracting hand-crafted features such as LBP maps [5], color components [6], and DCT maps [7] as informative indicators for Deepfake detection. Subsequent works focused on capturing anomalies like eye-blinking [8], head pose artifacts [9], and face-warping features [10]. Many methods aimed to improve detection accuracy by developing learning-based models using generic architectures such as Xception Net [11], Efficient Net [12], and Capsule Net [13]. Later approaches, including Face X-ray [14], SBI [15], DCL [16], and RECCE [16] sought to capture common forgery features to enhance model generalizability. With the advent of ViT [17], numerous ViT-based detectors [18, 19, 20, 21] have been developed. These approaches exploit global receptive fields and exhibit superior detection performance. However, previous face forgery detection methods still face two major challenges in real-world applications: (1) Existing methods struggle to generalize in uncontrolled environments. These models may encounter unforeseen forgery domains in the wild, which can lead to significant performance degradation; (2) In real-world applications, detection models typically need fine-tuning with newly acquired data to adapt to different domains. However, this adaptation process for pre-trained detection models can be computationally expensive, making it challenging to deploy them effectively, particularly on mobile devices with limited computational resources.
@@
Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from ImageNet weights. However, fully fine-tuning a network is costly, hindering its real-world applications. To mitigate this issue, we integrate lightweight Adapter and LoRA layers into the ViT backbone. During training, we freeze plain ViT backbones with ImageNet weights and solely update the lightweight Adapter and LoRA layers. These designed modules capture local and global forgery artifacts using slim components. Additionally, these modules preserve substantial ImageNet knowledge and enable flexible model adaptation for face forgery detection. While parameter-efficient fine-tuning (PEFT) has been widely adopted in various domains, our work makes a notable contribution by integrating PEFT with a forgery-aware perspective, specifically tailored for open-set face forgery detection—a task that presents unique challenges in generalizability and adaptation. Fig. 3 illustrates the average cross-dataset AUC performance across six unforeseen Deepfake datasets versus the number of activated parameters. This demonstrates that our method achieves superior generalizability with the fewest activated parameters.
@@
  • We design a forgery-style mixture module to enhance the diversity of source domains. The model’s generalizability is effectively improved by randomly mixing forgery styles.
@@
  • Our designed components jointly contribute to an effective and scalable solution for open-set Deepfake detection. Extensive experiments demonstrate that our model achieves state-of-the-art generalizability and robustness with only 1.34M activated parameters.
@@ II-A Face forgery detection
Early face forgery detection methods initially focused on extracting hand-crafted features, such as eye-blinking frequency [8], head pose inconsistency [9], face warping artifacts [10], and rPPG features [28]. With the rise of deep learning and artificial intelligence, numerous learning-based methods emerged to improve detection accuracy. Some generic CNNs, such as the Xception network [11, 29, 30, 7, 31], EfficientNet-b4 [12, 32, 33, 34], and Capsule network [13], have been utilized for face forgery detection, showing promising results. However, these methods often need to be more balanced in training data, leading to poor generalization capabilities. In response, some approaches have focused on extracting common forgery features. For example, Face X-ray [14] and SBI [15] aimed to extract common face blending features, enhancing model generalizability. F3Net and Luo et al. [30] mined frequency artifacts, while DCL [16] and CFM [32] employed contrastive learning schemes for more generalized Deepfake detection. NoiseDF [35] raised an innovative idea to distinguish Deepface-specific noise traces from common noises within each single frame, providing additional interpretability. With the rapid advancements in vision transformers (ViT), several ViT-based approaches [21, 18, 27, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] have been developed, showing superior detection performance. ICT [21] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection. However, the scalability of ViT-based methods is largely constrained by the quadratic complexity of attention computation, limiting their deployment, particularly on mobile devices with limited computing resources. In response, this paper introduces a parameter-efficient fine-tuning (PEFT) strategy for face forgery detection to address this issue. Furthermore, we combine different forgery styles to augment the forgery source domains, enhancing the model’s generalization capability.
@@ II-B Parameter-efficient fine-tuning (PEFT)
The attention mechanism has demonstrated significant success across various computer vision tasks. However, the self-attention in ViT scales quadratically concerning the input patch number, leading to substantial computational demands during training. Several methods have been proposed to address this issue, including parameter-efficient fine-tuning (PEFT). PEFT selectively updates a small portion of the model parameters, facilitating model adaptation to downstream tasks. Adapters [46, 47] are typically inserted between feedforward and normalization layers of transformers, comprising a down-sample and up-sample layer. For example, Low-Rank Adaptation (LoRA) [48] decomposes a matrix into two trainable low-rank matrices, significantly reducing the number of parameters compared to the original matrix. Visual Prompt Tuning (VPT) [49] introduces a small number of trainable parameters in the input space, treating additional tokens as learnable pixels. Convpass [50] aims to inject local priors into expressive ViT models by integrating convolutional operations, enhancing performance across various computer vision tasks. Neural Prompt Search (NOAH) [51] further advances by incorporating Adapter, LoRA, and VPT into vision transformers and optimizing their design through a neural architecture search algorithm. Despite these advancements, the development of PEFT methods for face forgery detection has been largely understudied, yet it is crucial in practical scenarios. In this paper, we integrate lightweight Adapter and LoRA modules into the plain ViT backbone, optimizing only these modules while keeping the ViT backbone fixed with ImageNet weights during training. The designed modules effectively capture global and local face forgery features more efficiently, without the danger of catastrophic forgetting typical of fully finetuning models [52].
@@ III-A Overall framework
Figure 6: Illustration of forgery style mixture pipeline. The designed feature extractor processes real faces and forgery source domain data, and the extracted features are subsequently delivered to the forgery-style mixture module.
@@ III-B Forgery-aware PEFT
where $x_{c}$ indicates the center pixel of the window $\Omega$, and $x_{5}$ represents $x_{c}$ in Fig. 5. $x_{i}$ indicates the peripheral pixels within $\Omega$. $w_{i}$ denotes the corresponding convolutional weights. Finally, the second $1\times 1$ Conv layer restores the channels of the feature, and the produced feature is flattened to the same shape as $h^{in}$. Our designed adapter layer effectively introduces inductive bias into standard ViTs, guiding the model in learning local artifacts for face forgery detection. Additionally, considering the effectiveness of high-frequency features in the task of face forgery detection [7, 53, 54, 55], our CDC operator enables the model to extract abundant high-frequency information from input faces, thereby enhancing detection performance.
@@
Figure 7: Pipeline of the proposed forgery feature mixture process.
The $W_{q/k/v}$ parameter denotes the pre-trained ImageNet weights fixed during the training process. Moreover, the LoRA module consists of two low-rank matrices:
@@
The designed forgery-aware PEFT modules bear the following advantages for face forgery detection: (1) By optimizing a small subset of parameters, our method reduces computational and memory requirements compared to full fine-tuning while also accelerating training, making frequent updates more practical in open-set environments. Moreover, storing the backbone weights in the cloud while keeping only a minimal set of parameters on edge devices facilitates deployment on resource-constrained mobile platforms. (2) As model sizes continue to grow, our forgery-aware PEFT strategy remains scalable for large models, and the designed modules can be seamlessly integrated with various backbones in a plug-and-play fashion. (3) Our approach efficiently incorporates forgery-specific knowledge while preserving the pretrained model’s general knowledge, effectively mitigating overfitting to forgery data and enhancing generalization performance.
@@ III-C Forgery style mixture
In face forgery detection, the training dataset often encompasses multiple forgery techniques. This training setting accommodates practical scenarios where available source domains are utilized. In this vein, we denote the training dataset as $D^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$. $R^{S}$ represents the real source faces, while $F_{i}^{S}$ denotes the $i_{th}$ source forgery domain, where $i\in\{1,2,...,N\}$. Following previous methods in face forgery detection, we use FF++ as the training set, which includes multiple training forgery domains (i.e., manipulation types). As such, the target testing domain can be denoted as $D^{T}=\{R^{T},F^{T}\}$. Similarly, $R^{T}$ and $F^{T}$ are target real and forgery domains. Our preliminary experiments, illustrated in Fig. 1 and Fig. 2, reveals a high false negative rate (FNR) and a low false positive rate (FPR) in cross-domain evaluations, suggesting a negligible domain gap between $R^{S}$ and $R^{T}$ and significant domain gaps between $F^{S}$ and $F^{T}$. As such, we propose to augment source forgery domains to mitigate this issue. Fig. 6 depicts the proposed forgery-style mixture pipeline. Each training batch, composed of one real domain and multiple fake domains, is processed by our proposed feature extractor, which shares weights across all source domains. The extracted features from the source forgery domains are passed to the designed forgery-style mixture module, and the mixed features are subsequently forwarded to the classifier to perform forgery detection.
@@
Finally, we restore the original index order of the mixed features $\mathcal{F}_{mix^{\prime}}$ to obtain $\mathcal{F}_{mix}$, which is forwarded to the classifier for final decision-making. Similar to [57], we set a default probability of 0.5 to activate the forgery-style mixture module. In the inference stage, the forgery-style mixture is deactivated. As such, the designed module augments the source forgery domains in the training phase, thereby mitigating the overfitting problem and enhance the model’s generalizability. This design effectively improves the model’s ability to generalize in open-set face forgery detection.
TABLE I: High quality (c23) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
@@
OSDFD Ours 1.34M 0.962 0.885 0.932 0.846 0.698 0.595 0.781 0.682 0.843 0.752
TABLE II: Low quality (c40) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
@@
OSDFD Ours 1.34M 0.853 0.762 0.717 0.663 0.713 0.661 0.630 0.595 0.728 0.670
@@ III-D Objective function
where $\lambda$ balances the two loss components. $L_{BCE}$ represents the binary cross-entropy loss. The Single-Center Loss, denoted as $L_{SCL}$, aims to compact the feature distribution of real faces while simultaneously moving fake features away from the center of real features:
@@
where $Dist_{R}$ and $Dist_{F}$ denote the average distance between the real feature center $C$ and real feature $f_{R}$ and fake feature $f_{F}$, respectively:
@@
We select the features after the second-to-last fully-connected layer to calculate $L_{SCL}$. This loss enhances the discriminative and separable characteristics of the real and fake face features, thereby improving the overall performance of face forgery detection.

### v3

@@ Abstract
Open-set face forgery detection poses significant security threats and presents substantial challenges for existing detection models. These detectors primarily have two limitations: they cannot generalize across unknown forgery domains or inefficiently adapt to new data. To address these issues, we introduce an approach that is both general and parameter-efficient for face forgery detection. Our method builds on the assumption that different forgery source domains exhibit distinct style statistics. Specifically, we design a forgery-style-mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains. In addition, previous methods typically require fully fine-tuning pretrained networks, consuming substantial time and computational resources. Drawing on recent advancements in vision transformers (ViT) for face forgery detection, we develop a parameter-efficient ViT-based detection model that includes lightweight forgery feature extraction modules and enables the model to extract global and local forgery clues simultaneously. We only optimize the inserted lightweight modules during training, maintaining the original ViT structure with its pre-trained weights. This training strategy effectively preserves the informative pre-trained knowledge while flexibly adapting the model to the task of Deepfake detection. Extensive experimental results demonstrate that the designed model achieves state-of-the-art generalizability with significantly reduced trainable parameters, representing an important step toward open-set Deepfake detection in the wild.
@@ I Introduction
Face forgery detection has long posed a challenging yet crucial problem. Early methodologies proposed extracting hand-crafted features such as LBP maps [5], color components [6], and DCT maps [7] as informative indicators for Deepfake detection. Subsequent works focused on capturing anomalies like eye-blinking [8], head pose artifacts [9], and face-warping features [10]. Many methods aimed to improve detection accuracy by developing learning-based models using generic architectures such as Xception Net [11], Efficient Net [12], and Capsule Net [13]. Later approaches, including Face X-ray [14], SBI [15], DCL [16], and RECCE [16] sought to capture common forgery features to enhance model generalizability. With the advent of ViT [17], numerous ViT-based detectors [18, 19, 20, 21, 22] have been developed. These approaches exploit global receptive fields and exhibit superior detection performance. However, previous face forgery detection methods still face two major challenges in real-world applications: (1) Existing methods struggle to generalize in uncontrolled environments. These models may encounter unforeseen forgery domains in the wild, which can lead to significant performance degradation; (2) In real-world applications, detection models typically need fine-tuning with newly acquired data to adapt to different domains. However, this adaptation process for pre-trained detection models can be computationally expensive, making it challenging to deploy them effectively, particularly on mobile devices with limited computational resources.
@@
Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from weights. However, fully fine-tuning a network is costly, hindering its real-world applications. To mitigate this issue, we integrate lightweight Adapter and LoRA layers into the ViT backbone. During training, we freeze plain ViT backbones with weights and solely update the lightweight Adapter and LoRA layers. These designed modules capture local and global forgery artifacts using slim components. Additionally, these modules preserve substantial ImageNet knowledge and enable flexible model adaptation for face forgery detection. While parameter-efficient fine-tuning (PEFT) has been widely adopted in various domains, our work makes a contribution by integrating PEFT with a forgery-aware perspective, specifically tailored for open-set face forgery detection—a task that presents unique challenges in generalizability and adaptation. Fig. 3 illustrates the average cross-dataset AUC performance across six unforeseen Deepfake datasets versus the number of activated parameters. This demonstrates that our method achieves superior generalizability with the fewest activated parameters.
@@
  • We design a forgery-style-mixture module to enhance the diversity of source domains. The model’s generalizability is effectively improved by randomly mixing forgery styles.
@@
  • Our designed components jointly contribute to an effective and scalable solution for open-set Deepfake detection. Extensive experiments demonstrate that our model achieves state-of-the-art generalizability and robustness with the fewest activated parameters.
@@ II-A Face forgery detection
Early face forgery detection methods initially focused on extracting hand-crafted features, such as eye-blinking frequency [8], head pose inconsistency [9], face warping artifacts [10], and rPPG features [29]. With the rise of deep learning and artificial intelligence, numerous learning-based methods emerged to improve detection accuracy. Some generic CNNs, such as the Xception network [11, 30, 31, 7, 32], EfficientNet-b4 [12, 33, 34, 35, 36], and Capsule network [13], have been utilized for face forgery detection, showing promising results. However, these methods often need to be more balanced in training data, leading to poor generalization capabilities. In response, some approaches have focused on extracting common forgery features. For example, Face X-ray [14] and SBI [15] aimed to extract common face blending features, enhancing model generalizability. F3Net and Luo et al. [31] mined frequency artifacts, while DCL [16] and CFM [33] employed contrastive learning schemes for more generalized Deepfake detection. NoiseDF [37] raised an innovative idea to distinguish Deepface-specific noise traces from common noises within each single frame, providing additional interpretability. With the rapid advancements in vision transformers (ViT), several ViT-based approaches [21, 18, 28, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 39] have been developed, showing superior detection performance. ICT [21] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans [48] designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection. Forensics Adapter [49] is a lightweight network that equips CLIP with forgery-specific knowledge by learning blending boundary traces and enhancing visual tokens through a dedicated interaction strategy. TFCU [50] improves face forgery video detection by sequentially capturing momentary anomalies, gradual inconsistencies, and cumulative distortions through its consecutive correlation, future-guided, and historical review modules. OWG-DS [51] advances real-world Deepfake detection by transferring knowledge from limited labeled data to large-scale unlabeled data. GM-DF [52] enhances Deepfake detection generalization across multiple datasets, enabling effective handling of diverse real-world scenarios and unseen attacks. DAID [53] mitigates demographic biases in Deepfake detection through demographic-aware data rebalancing and demographic-agnostic feature aggregation. Recent LVLM-based approaches [54, 55, 56, 57, 58, 59, 60] take a further step forward by achieving explainable face forgery detection. However, the scalability of ViT-based methods is largely constrained by the quadratic complexity of attention computation, limiting their deployment, particularly on mobile devices with limited computing resources. In response, this paper introduces a parameter-efficient fine-tuning (PEFT) strategy for face forgery detection to address this issue. Furthermore, we combine different forgery styles to augment the forgery source domains, enhancing the model’s generalization capability.
@@ II-B Parameter-efficient fine-tuning (PEFT)
The attention mechanism has demonstrated significant success across various computer vision tasks. However, the self-attention in ViT scales quadratically concerning the input patch number, leading to substantial computational demands during training. Several methods have been proposed to address this issue, including parameter-efficient fine-tuning (PEFT). PEFT selectively updates a small portion of the model parameters, facilitating model adaptation to downstream tasks. Adapters [61, 62] are typically inserted between feedforward and normalization layers of transformers, comprising a down-sample and up-sample layer. For example, Low-Rank Adaptation (LoRA) [63] decomposes a matrix into two trainable low-rank matrices, significantly reducing the number of parameters compared to the original matrix. Visual Prompt Tuning (VPT) [64] introduces a small number of trainable parameters in the input space, treating additional tokens as learnable pixels. Convpass [65] aims to inject local priors into expressive ViT models by integrating convolutional operations, enhancing performance across various computer vision tasks. Neural Prompt Search (NOAH) [66] further advances by incorporating Adapter, LoRA, and VPT into vision transformers and optimizing their design through a neural architecture search algorithm. Despite these advancements, the development of PEFT methods for face forgery detection has been largely understudied, yet it is crucial in practical scenarios. In this paper, we integrate lightweight Adapter and LoRA modules into the plain ViT backbone, optimizing only these modules while keeping the ViT backbone fixed with pretrained weights during training. The designed modules effectively capture global and local face forgery features more efficiently, without the danger of catastrophic forgetting typical of fully finetuning models [67].
@@ III-B Forgery-aware PEFT
where $x_{c}$ indicates the center pixel of the window $\Omega$, and $x_{5}$ represents $x_{c}$ in Fig. 5. $x_{i}$ indicates the peripheral pixels within $\Omega$. $w_{i}$ denotes the corresponding convolutional weights. Finally, the second $1\times 1$ Conv layer restores the channels of the feature, and the produced feature is flattened to the same shape as $h^{in}$. CDC operates within a local sliding window, calculating differences between peripheral and central pixels to extract high-frequency features [68]. The CDC adapter functions as a local anomaly extractor for the ViT backbone by capturing subtle forgery cues, such as boundary inconsistencies and local irregularities. These local cues are then integrated with the backbone’s global features, while the PEFT strategy ensures this integration is achieved effectively with minimal trainable parameters. Our designed adapter layer effectively introduces inductive bias into standard ViTs, guiding the model in learning local artifacts for face forgery detection. Additionally, considering the effectiveness of high-frequency features in the task of face forgery detection [7, 69, 70, 71], our CDC operator enables the model to extract abundant high-frequency information from input faces, thereby enhancing detection performance.
@@
Figure 6: Illustration of forgery style mixture pipeline. The designed feature extractor processes real faces and forgery source domain data, and the extracted features are subsequently delivered to the forgery-style-mixture module.
@@
The $W_{q/k/v}$ parameter denotes the pre-trained weights fixed during the training process. Moreover, the LoRA module consists of two low-rank matrices:
@@
The designed forgery-aware PEFT modules bear the following advantages for face forgery detection: (1) By optimizing a small subset of parameters, our method reduces computational and memory requirements compared to full fine-tuning while also accelerating training, making frequent updates more practical in open-set environments. Moreover, storing the backbone weights in the cloud while keeping only a minimal set of parameters on edge devices facilitates deployment on resource-constrained mobile platforms. (2) As model sizes continue to grow, our forgery-aware PEFT strategy remains scalable for large models, and the designed modules can be seamlessly integrated with various transformer-based backbones in a plug-and-play fashion. (3) Our approach efficiently incorporates forgery-specific knowledge while preserving the pretrained model’s general knowledge, effectively mitigating overfitting to forgery data and enhancing generalization performance.
Figure 7: Pipeline of the proposed forgery feature mixture process.
@@ III-C Forgery style mixture
In face forgery detection, the training dataset often encompasses multiple forgery techniques. This training setting accommodates practical scenarios where available source domains are utilized. In this vein, we denote the training dataset as $D^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$. $R^{S}$ represents the real source faces, while $F_{i}^{S}$ denotes the $i_{th}$ source forgery domain, where $i\in\{1,2,...,N\}$. Following previous methods in face forgery detection, we use FF++ as the training set, which includes multiple training forgery domains (i.e., manipulation types). As such, the target testing domain can be denoted as $D^{T}=\{R^{T},F^{T}\}$. Similarly, $R^{T}$ and $F^{T}$ are target real and forgery domains. Our preliminary experiments, illustrated in Fig. 1 and Fig. 2, reveals a high false negative rate (FNR) and a low false positive rate (FPR) in cross-domain evaluations, suggesting a negligible domain gap between $R^{S}$ and $R^{T}$ and significant domain gaps between $F^{S}$ and $F^{T}$. As such, we propose to augment source forgery domains to mitigate this issue. Fig. 6 depicts the proposed forgery-style-mixture pipeline. Each training batch, composed of one real domain and multiple fake domains, is processed by our proposed feature extractor, which shares weights across all source domains. The extracted features from the source forgery domains are passed to the designed forgery-style-mixture module, and the mixed features are subsequently forwarded to the classifier to perform forgery detection.
@@
Finally, we restore the original index order of the mixed features $\mathcal{F}_{mix^{\prime}}$ to obtain $\mathcal{F}_{mix}$, which is forwarded to the classifier for final decision-making. Similar to [73], we set a default probability of 0.5 to activate the forgery-style-mixture module. In the inference stage, the forgery-style-mixture is deactivated. As such, the designed module augments the source forgery domains in the training phase, thereby mitigating the overfitting problem and enhance the model’s generalizability. This design effectively improves the model’s ability to generalize in open-set face forgery detection.
In contrast to conventional computer vision tasks, forgery detection suffers from imbalanced domain gaps: the discrepancies between different fake domains are substantially larger than those among real domains, due to the continual evolution of manipulation techniques. Forgery Style Mixture (FSM) is a label-guided mechanism that explicitly mixes feature statistics across different forgery styles. Unlike prior augmentation strategies that enhance general feature diversity, FSM specifically targets forgery-style diversity, thereby expanding the fake feature space without requiring additional data.
TABLE I: High quality (c23) cross-manipulation detection performance. (#Params: trainable parameter number)
@@
CLIP [83] ICML21 303.18M 0.890 0.766 0.847 0.737 0.541 0.504 0.652 0.538 0.733 0.636
OSDFD (ViT-B) Ours 1.34M 0.962 0.885 0.932 0.846 0.698 0.595 0.781 0.682 0.843 0.752
OSDFD (CLIP) Ours 2.89M 0.990 0.934 0.945 0.858 0.881 0.790 0.793 0.703 0.902 0.821
TABLE II: Low quality (c40) cross-manipulation detection performance. (#Params: trainable parameter number)
@@
CLIP [83] ICML21 303.18M 0.800 0.707 0.676 0.615 0.651 0.606 0.592 0.544 0.680 0.618
OSDFD (ViT-B) Ours 1.34M 0.853 0.762 0.717 0.663 0.713 0.661 0.630 0.595 0.728 0.670
OSDFD (CLIP) Ours 2.89M 0.924 0.819 0.761 0.684 0.817 0.716 0.628 0.601 0.783 0.705
@@ III-D Objective function
$\lambda$ balances the two loss components. $L_{BCE}$ represents the binary cross-entropy loss. The Single-Center Loss, denoted as $L_{SCL}$, aims to compact the feature distribution of real faces while simultaneously moving fake features away from the center of real features:
@@
where margin (default: 0.01) is a hyperparameter that specifies how much farther the fake features should be from the real center compared to the real features. $Dist_{R}$ and $Dist_{F}$ denote the average distance between the real feature center $C$ and real feature $f_{R}$ and fake feature $f_{F}$, respectively:
@@
We select the features after the second-to-last fully-connected layer to calculate $L_{SCL}$. This loss enhances the discriminative and separable characteristics of the real and fake face features, thereby improving the overall performance of face forgery detection. FSM is coupled with Single Center Loss ($L_{SCL}$) to jointly form a forgery-oriented regularization strategy.